OpenStudy (anonymous):

Hunter placed 2 red, 3 blue, 1 yellow, 8 black, and 7 white straws into a bag. What is the probability that he will pull out a white straw when he reaches into the bag?

hartnn (hartnn):

how many white straws are there and how many total straws are there ?

OpenStudy (anonymous):

7 white and 21 total

OpenStudy (anonymous):

do i divide that

hartnn (hartnn):

yup.

OpenStudy (lstrasmann):

so divide 21 by 7

OpenStudy (anonymous):

cause i got three

OpenStudy (anonymous):

so 1/3

hartnn (hartnn):

the probability will be no. of white straws / total yes 1/3 :)

OpenStudy (lstrasmann):

oh yeah 7 divided by 21 not 21 divided by 7

OpenStudy (anonymous):

so if i am doing a similar problem to that and the total is 19 and i am dividing that by 4 to get 4.75 what does that become when i am needing to change it to a fraction?

OpenStudy (anonymous):

4/19? 7/19?

OpenStudy (lstrasmann):

you divide the number of *whatever* by the total number, so 4 divided by 19 is 0.21 which simplifies to\[\frac{ 21 }{ 100 } \]

hartnn (hartnn):

yup, 4/19

OpenStudy (anonymous):

well that cant be right because the different answer choices i have are 4/19 7/19 5/19 3/19

hartnn (hartnn):

4/19 is one of them.

OpenStudy (anonymous):

thanks hartnn

OpenStudy (lstrasmann):

oh yeah forgot that fractions are just division :) lol don't listen to me

OpenStudy (anonymous):

i meant that towards istrasmann

OpenStudy (lstrasmann):

its Lstrasmann

OpenStudy (anonymous):

;p lol

OpenStudy (anonymous):

Marty has a standard deck containing 52 cards. If Marty takes one card from the deck, what is the probability that he will select a king of hearts?

OpenStudy (anonymous):

what is it as easy as 1/52

hartnn (hartnn):

yup. 1/52

hartnn (hartnn):

because there is only 1 king of heart.

OpenStudy (anonymous):

oh didnt know thats why i wasnt sure lol so back to the first type of questions i was asking i understand how to do the problem just not how to convert to fractions

OpenStudy (anonymous):

so if the total is 15 and i am dividing by 3 which equals 5 how do i make it a fraction

OpenStudy (lstrasmann):

you divide 3 by 15, like this: \[\frac{ 3 }{ 15 }\]

hartnn (hartnn):

TOTAL is always in denominator! for probability problems... so it would be 3/15

hartnn (hartnn):

Probability = Favorable number of events / Total number of events.

OpenStudy (anonymous):

well thats what i have been assuming but the answers i can choose from are different 1/5 1/15 3 2/15 so would it be 3 in this case?

OpenStudy (lstrasmann):

3/15 simplifies to 1/5 so its 1/5

OpenStudy (anonymous):

so how did you figure that out just curious

OpenStudy (anonymous):

did you divide that problem by 3 to simplify it since 3 is 1 as 15 is to 5

OpenStudy (lstrasmann):

\[\frac{ 3 }{ 15} \div \frac{ 3 }{ 3 } = \frac{ 1 }{ 5 }\]

OpenStudy (lstrasmann):

since 3/3 is basically 1, so it would be like dividing 3 by 3 to get 1 and 15 by 3 to get 5

OpenStudy (anonymous):

so if i am doing another one with the total 20 and dividing it by 4 would i get 1/4?

OpenStudy (lstrasmann):

no 1/5

hartnn (hartnn):

4/20=1/5

OpenStudy (anonymous):

6/14 1/4 5/19 3/20 possible answers this is why im confused because every set of answers is different from what i think it should be like 1/5 ya know its weird and confusing ugh

OpenStudy (lstrasmann):

maybe you totaled them wrong.. try totaling them up again

hartnn (hartnn):

or maybe u counted 4 as wrong...

OpenStudy (anonymous):

3+4+2+5+6= 20 divided by 4 so i dont know what i did wrong

OpenStudy (anonymous):

oh god it was not 4 it was 5 divided by 5 pellet lol

hartnn (hartnn):

5/20 = 1/4

OpenStudy (lstrasmann):

ok then its 1/4

hartnn (hartnn):

Remember to always keep the total in the denominator.

OpenStudy (anonymous):

awesome thanks for now got to go eat lunch haha

OpenStudy (lstrasmann):

you're welcome :) have a great lunch